OpenStudy (anonymous):

multiply (2 sqrt 5 - 1)(3 sqrt 5+4)

OpenStudy (anonymous):

\[\LARGE Multiply (2\sqrt{5}-1)(3\sqrt{5}+4)\]

OpenStudy (anonymous):

30-5 sqrt 5.

OpenStudy (anonymous):

sorry. answer is 30+5 sqrt 5 :)

OpenStudy (anonymous):

That isn't one of the options. Here they are: \[\Large A) 26-11\sqrt{5}\] \[\Large B) 26+5\sqrt{5}\] \[\Large C) 34-11\sqrt{5}\] \[\Large D) 34+5\sqrt{5}\]

OpenStudy (anonymous):

sorry again. answer is 26+5 sqrt 5 :) :) :)

OpenStudy (anonymous):

are you sure? Lol

OpenStudy (anonymous):

yes i am.

OpenStudy (anonymous):

Thanks! :)

OpenStudy (anonymous):

welcome. i guess i need a rest :)